# I Know What I Don’t Know: Improving Model Cascades Through Confidence Tuning

### v1

@@ 4   Experiments
In this section we detail the experiments used to evaluate our deferral strategies across three distinct model architectures: encoder-only classification models, decoder-only LMs, and encoder-decoder VLMs. Each setup leverages a deferral setup from smaller to larger models, assessing the efficacy of deferring hard queries to more capable models.
@@ 4.1 Encoder-only Setup (Classification Models)
We comprehensively assess the performance of our deferral strategies across different model architectures and task types, starting with image classification. We train both a large model and a small model on the following datasets: CIFAR-10/100 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and TinyImageNet200 (Le & Yang, 2015). For both CIFAR datasets we use a ResNet-18 (He et al., 2016) as $\mathcal{M}_{L}$ and a custom CNN as $\mathcal{M}_{S}$. For Food-101 and TinyImageNet200 we instead use a ResNet-50 (He et al., 2016) as $\mathcal{M}_{L}$ and a Mobilenet V3 Small (Howard et al., 2019) as $\mathcal{M}_{S}$, where the latter is trained using knowledge distillation from the big model.
Evaluation Metrics.
Figure 4: Performance on image classification tasks. We observe that lower levels of $\alpha$ lead to decreased distributional overlap between correct/incorrect predictions (left), increased deferral performance (center) and generally decreased performance over the full data distribution (right). These results support our conclusion that the small model $\mathcal{M}_{S}$ learns to refocus on easier subsets of the distribution while understanding more reliably when it should abstain and defer to the large model $\mathcal{M}_{L}$.
We measure the performance of our confidence tuning method and the resulting deferral function $g(\cdot)$ using the following performance metrics (see example in Figure 3 for a visual overview):
@@
    The Distributional Overlap of Confidences of Correct and Incorrect Predictions $s_{o}$ is defined as the integral of the minimum of the probability density functions (PDFs) of confidence scores for correctly classified samples, $\hat{p}_{\text{corr}}(c)$, and incorrectly classified samples, $\hat{p}_{\text{incorr}}(c)$ (see Figure 3a). Formally, given the confidence sets $\mathcal{C}_{\text{corr}}$ and $\mathcal{C}_{\text{incorr}}$, the overlap $s_{o}$ is computed as
    $s_{o}=\int_{0}^{1}\min\left\{\hat{p}_{\text{corr}}(c),\ \hat{p}_{\text{incorr}}(c)\right\}\,\mathrm{d}c,$ (9)
    where the PDFs are estimated using Kernel Density Estimation (KDE). If $s_{o}=1$, then $\mathcal{M}_{S}$ cannot distinguish the confidence distribution of correct and incorrect predictions; if $s_{o}=0$, then $\mathcal{M}_{S}$ can perfectly separate correct and incorrect predictions. Note that a related way of capturing the distributional separability is given by the Area Under the Receiver Operating Characteristic Curve (AUROC) which we discuss in Appendix B.3.
@@
    Deferral Performance $s_{d}$: To formally quantify how well $\mathcal{M}_{S}$ defers difficult inputs to $\mathcal{M}_{L}$, we examine the joint performance across all possible deferral ratios $r\in[0,1]$, where $r$ denotes the fraction of inputs sent to $\mathcal{M}_{L}$ based on a particular threshold $\tau$ (recall Equation (6)). Figure 3 b) illustrates how, as $r$ increases from 00 to $1$, the overall (joint) accuracy $\text{acc}(r)$ increases from the accuracy of $\mathcal{M}_{S}$ (blue circle, no deferral) to the accuracy of $\mathcal{M}_{L}$ (orange square, full deferral). Useful deferral models are constrained to operate between random deferral ($\mathrm{acc}_{\mathrm{rand}}$, red line) and ideal deferral ($\mathrm{acc}_{\mathrm{ideal}}$, green line). The ideal deferral $\mathrm{acc}_{\mathrm{ideal}}$ corresponds to the oracle solution that perfectly defers examples misclassified by $\mathcal{M}_{S}$ and we discuss its exact functional form in Appendix A.2. We also define the realized deferral curve, $\mathrm{acc}_{\mathrm{real}}$, as the joint accuracy obtained under the learned deferral strategy $g(\cdot)$ employed by $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$. The deferral performance metric $s_{d}$ is then given as:
    $\small s_{d}=\frac{A_{\mathrm{perf}}}{A_{\mathrm{useful}}}=\frac{\int_{0}^{1}\left(\mathrm{acc}_{\mathrm{real}}(r)-\mathrm{acc}_{\mathrm{rand}}(r)\right)\,\mathrm{d}r}{\int_{0}^{1}\left(\mathrm{acc}_{\mathrm{ideal}}(r)-\mathrm{acc}_{\mathrm{rand}}(r)\right)\,\mathrm{d}r}.$ (10)
    This ratio quantifies the fraction of the potential improvement over random deferral that has been realized by the achieved deferral strategy. Note that $s_{d}=1$ indicates perfect deferral, matching the ideal strategy, while an $s_{d}=0$ implies no improvement over random deferral.
@@
    Accuracy of the small model $\text{acc}(\mathcal{M}_{S})$: Finally, since Gatekeeper emphasizes patterns for distinguishing correct/incorrect examples, the model is no longer encouraged to minimize the classification loss over the full population. As a result, improving on the correct/incorrect separation task can lead to drops in utility. Hence, practically useful deferral methods need to balance both deferral performance and the accuracy of $\mathcal{M}_{S}$.
Figure 6: Performance on language modeling tasks. Similar as Figure 4. In addition to a non-tuned baseline, we also add an uncertainty prompting baseline as well as an Answer “N” option. Both these approaches fail to meaningfully improve deferral.
Results. We document our main results in Figure 4. We report performance results on both a baseline model, i.e., an instance of $\mathcal{M}_{S}$ that was not trained with Gatekeeper, and a set of small models trained with Gatekeeper at various $\alpha$s. Across all models we compute the deferral performance as well as the model’s correct/incorrect separation ability (center, left). We see that the strongest performance is achieved at low $\alpha$s as the model strongly emphasizes pushing the outputs of incorrect examples closer to a uniform distribution. However, this strong performance comes at a cost: the accuracy of the small model consistently degrades for small $\alpha$s (right). This highlights the fact that the model effectively “unlearns” to perform well on some part of the data distribution and re-focuses its classification ability of easier data points. On the other hand, for large values of $\alpha$ we observe consistency in model accuracy or even slight improvements as the model now emphasizes most of its training on maintaining good performance on already well-predicted points.
This result highlights a critical trade-off which is directly controlled by $\alpha$: how strongly do we want to degrade model performance over the full data distribution in order to obtain a better deferral model? We note that this compromise between raw model utility and deferral performance is not surprising and similar trade-offs exist in fairness (Dutta et al., 2020; Yaghini et al., 2023) and privacy (Abadi et al., 2016; Rabanser et al., 2023). We study this trade-off explicitly in Figure 5 showing (i) a clear negative correlation between deferral performance and the small model’s accuracy; and (ii) a clear positive correlation between the overlap of correct/incorrect confidences and the accuracy of $\mathcal{M}_{S}$.
@@ 4.2 Decoder-only Setup (Language Models)
In the decoder-only setup, we explore the application of LLMs. Our primary models of interest are the scalable LMs from the Gemma model class (GemmaTeam et al., 2024). We choose Gemma2B as $\mathcal{M}_{S}$ and Gemma7B as $\mathcal{M}_{L}$ with 2 billion and 7 billion parameters, respectively. Similar to the encoder-only setup, we employ smaller LMs as the initial classifiers to manage simpler next-token prediction tasks. The deferral strategy involves routing only those token sequences that exhibit high uncertainty — as determined by high predictive entropy — to the more powerful model $\mathcal{M}_{L}$.
Our experiments start by taking the instruction-tuned checkpoints of Gemma2B and Gemma7b and fine-tuning both models on the training split of a respective dataset to ensure that the model (i) performs well on the task; and (ii) is familiar with the desired response format. Note that this fine-tuning stage is performed using standard perplexity minimization. Then, we finetune $\mathcal{M}_{S}$ with Gatekeeper using the same training split to decrease confidence on incorrect next-token predictions. Finally, we evaluate the model yielded by Gatekeeper on a validation/test split. The datasets we consider are ARC-e/c (Clark et al., 2018), MMLU (Hendrycks et al., 2020), and GSM8K (Cobbe et al., 2021). We use the same evaluation metrics as previously used in Section 4.1.
Results. We document our main results in Figure 6 where we compare the baseline model’s deferral and correct/incorrect separation ability against our fine-tuned model at different $\alpha$s. We generally observe a similar trend as in the image classification results: higher $\alpha$s maintain raw prediction performance closer to the baseline model but do not significantly improve correct/incorrect separation. At the same time, low $\alpha$s improve deferral more substantially at the cost of accuracy on the full data distribution. In addition to the baseline model (i.e., a model that was not fine-tuned with Gatekeeper but from which we still compute the predictive entropy as a deferral signal), we also include results for two other uncertainty prompting baselines (details in Appendix B.2): (i) Reduce Confidence: where we append additional instructions to the input prompt to encourage the model to reduce confidence when it is uncertain; and (ii) Answer “N”: where we instruct the model to answer with “N” if it is uncertain about the answer. Consistent with recent findings in Kadavath et al. (2022), we find that these approaches do not reliably improve separation of correct-incorrect predictions or offer advantages as deferral models.
Figure 7: Performance on VLM classification (left) and captioning tasks (right). Consistent with results in Figures 4 and 6, we see that smaller $\alpha$s lead to improved deferral performance in both classification and generation tasks.
@@ 4.3 Encoder-Decoder Setup (Vision-Language Models)
Finally, we examine models that incorporate both visual and textual processing capabilities, making it ideal for tasks that require a comprehensive understanding of image content in conjunction with language generation. We consider the PaliGemma (Steiner et al., 2024) model family which are encoder-decoder models designed to perform VL tasks such as image captioning, visual question answering, and image classification with descriptive outputs. In this setup, the encoder component processes the input images to extract rich feature representations, while the decoder generates corresponding textual classifications or descriptions. We use PaliGemma1B as $\mathcal{M}_{S}$ and PaliGemma7B as $\mathcal{M}_{L}$. The deferral strategy involves deploying a smaller VLM to handle the majority of classification tasks, reserving the more resource intensive 7B model for instances where $\mathcal{M}_{S}$’s predictive entropy falls below a predefined threshold.
Similarly to our experiments on LMs in Section 4.2, we employ two stages of fine-tuning. First, we take the instruction-tuned checkpoints of PaliGemma1B and PaliGemma7B and then fine-tune both models on the training split of a given dataset. Next, we fine-tune only $\mathcal{M}_{S}$ using Gatekeeper before evaluating the model on a validation/test split of the dataset. The datasets we consider are two classification datasets (VQAv2 (Goyal et al., 2017), AI2D (Hiippala et al., 2021)) and two captioning datasets (Cococap (Lin et al., 2014), Screen2Words (Wang et al., 2021)). This allows us to evaluate Gatekeeper in both closed-form vision-language classification setups as well as open-form text generation.
Factuality Scoring. For classification tasks we apply our analysis in the same way as in Section 4.2. However, for captioning datasets we need to evaluate the quality of a caption generated by PaliGemma. To do that, we compute a factuality score which judges whether the generated caption is semantically coherent with respect to a reference caption using the Gemini LLM (GeminiTeam et al., 2023). Specifically, the Gemini LLM is prompted with an instruction of the form: “Are these captions semantically equivalent?”, followed by both the candidate caption and the reference caption. The model then responds with either “Yes” or “No”. Finally, we compute the log-likelihood of each response and normalize it to a probability, reflecting the LLM’s confidence in the captions being factually aligned. We detail this process in Appendix B.4 and denote the factuality score for input point $\mathbf{x}_{i}$ with candidate caption $\hat{\mathbf{y}}_{i}$ and ground truth caption $\mathbf{y}_{i}$ as $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$.
Measuring Correlation Between Factuality and Negative Predictive Entropy. Since the result of evaluating $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$ is no longer binary, our evaluation metrics which previously relied on accuracy cannot be used directly to evaluate deferral performance and the correct/incorrect entropy distribution separation. We address this issue by replacing the distributional overlap computation with the Pearson correlation $\rho(g_{\text{NENT}}(\mathbf{x}_{i}),s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i}))$ between the negative predictive entropy of a caption $g_{\text{NENT}}(\mathbf{x}_{i})$ and its associated factuality score $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i}))$. We also adapt our deferral performance metric from Equation (10) to rely on factuality measures instead of accuracy.
Results. We document our results in Figure 7 where we compare the baseline model’s deferral ability against our fine-tuned models at different $\alpha$s. For the classification results (Figure 7 left), we observe the same trends as outlined in our classification and language modeling experiments. For the captioning results (Figure 7 right) we observe that Gatekeeper measurably increases the correlation between factuality and negative predictive entropy, yielding better deferral from $\mathcal{M}_{S}$ to $\mathcal{M}_{L}$ with decreasing $\alpha$. This demonstrates that our method does not just work on classification problems but also generalizes to sequence generation tasks. Note that while we also benchmarked the prompting baselines from Section 4.2 for these experiments, the PaliGemma model did not return any responses for these modified prompts (likely due to PaliGemma’s rigid pretraining and prompting instructions (Beyer et al., 2024)).
@@ 5   Conclusion
Limitations While our approach demonstrates promising results, there are a few notable constraints. First, we assume that only the smaller model can be tuned, while in some application the larger model might also adjustable for deferral. Second, our experiments primarily measure improvement over a single untuned baseline, potentially overlooking broader comparative insights. Third, we did not extensively evaluate across different model families in LLM and VLM settings (although we did so in classification experiments with ResNet vs. MobileNet). Finally, using a generative model (e.g., Gemini) to judge VLM captioning introduces the risk of erroneous assessments since LLMs are also imperfect oracles.
Appendix A Additional Background
A.1 Model Access Levels

### v2

@@ 4 Experiments
In this section, we detail the experiments used to evaluate the effectiveness of Gatekeeper across three distinct model architectures: encoder-only classification models, decoder-only language models, and encoder-decoder vision–language models. Each setup involves a cascade where a smaller model defers uncertain inputs to a larger model.
@@ 4.1 Encoder-only Setup (Classification Models)
We begin our evaluation with image classification using encoder-only models. We train a small model $\mathcal{M}_{S}$ and a larger model $\mathcal{M}_{L}$ on CIFAR-10/100 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and TinyImageNet200 (Le & Yang, 2015). For CIFAR, $\mathcal{M}_{L}$ is a ResNet-18 and $\mathcal{M}_{S}$ is a custom CNN. For Food-101 and TinyImageNet, $\mathcal{M}_{L}$ is a ResNet-50 and $\mathcal{M}_{S}$ is a MobileNet V3 Small (Howard et al., 2019) trained by distilling $\mathcal{M}_{L}$.
Figure 4: Performance on image classification tasks. We observe that lower levels of $\alpha$ lead to decreased distributional overlap between correct/incorrect predictions (left), increased deferral performance (center) and generally decreased performance over the full data distribution (right). These results support our conclusion that the small model $\mathcal{M}_{S}$ learns to refocus on easier subsets of the distribution while understanding more reliably when it should defer to the large model $\mathcal{M}_{L}$.
Evaluation Metrics. We evaluate the effectiveness of Gatekeeper and the resulting deferral function $g(\cdot)$ using three metrics that reflect different aspects of performance (see Figure 3 for an illustrative example):
@@
    Distributional Overlap of Confidence Scores $s_{o}$: Measures how well $\mathcal{M}_{S}$ separates correct from incorrect predictions based on output confidence. We estimate KDEs over confidence scores for correctly ($\hat{p}_{\text{corr}}$) and incorrectly ($\hat{p}_{\text{incorr}}$) classified samples and define:
[…]
    Lower $s_{o}$ indicates better separability ($s_{o}=0$ for perfect, $s_{o}=1$ for complete overlap). This metric is related to AUROC but compares full probability mass rather than prediction ranking.
@@
    Deferral Performance $s_{d}$: Quantifies how effectively $\mathcal{M}_{S}$ defers uncertain inputs to $\mathcal{M}_{L}$. We compute the joint accuracy $\text{acc}_{\text{real}}(r)$ over varying deferral rates $r$, and compare it against two baselines: random deferral $\text{acc}_{\text{rand}}(r)$ and ideal deferral $\text{acc}_{\text{ideal}}(r)$. The score is the normalized area between real and random deferral:
[…]
    A value of $s_{d}=1$ indicates optimal deferral; $s_{d}=0$ implies no improvement over random routing. See Appendix B.3 for details.
@@
    Small Model Accuracy $\text{acc}(\mathcal{M}_{S})$: Since Gatekeeper prioritizes uncertainty calibration over classification accuracy, $\mathcal{M}_{S}$ may sacrifice performance on the full data distribution to better recognize its errors. We report $\text{acc}(\mathcal{M}_{S})$ to quantify this trade-off and assess utility loss when optimizing for deferral.
Results. Figure 4 shows results for models trained with varying $\alpha$ values, compared against a baseline model and the cascading method of Narasimhan et al. (2022). At low $\alpha$, Gatekeeper effectively separates correct and incorrect confidence scores (lower $s_{o}$), improving deferral performance $s_{d}$. However, this comes with reduced small model accuracy, as $\mathcal{M}_{S}$ increasingly focuses on easy examples and assigns lower confidence to hard ones. As $\alpha$ increases, accuracy improves or stabilizes, but gains in deferral diminish. Notably, the baseline from Narasimhan et al. (2022) maintains model accuracy but requires auxiliary mechanisms to predict expert correctness—such as additional heads or networks—which can increase deployment complexity. In contrast, Gatekeeper is architecture-agnostic and operates purely via confidence calibration.
The accuracy-deferral trade-off is further explored in Figure 5, where we observe: (i) a negative correlation between deferral performance $s_{d}$ and small model accuracy, and (ii) a strong relationship between confidence overlap $s_{o}$ and both deferral and accuracy. The effect mirrors well-known trade-offs in fairness (Dutta et al., 2020; Yaghini et al., 2023) and privacy (Abadi et al., 2016; Rabanser et al., 2023). Crucially, Gatekeeper exposes this trade-off explicitly via a single tunable parameter $\alpha$, allowing practitioners to tailor cascaded systems to their computational or reliability needs.
Figure 6: Performance on language modeling tasks. Similar as Figure 4. In addition to a non-tuned baseline, we also add an uncertainty prompting baseline, an Answer “N” option, as well as the post-hoc confidence calibration method from Gupta et al. (2024). We observe that Gatekeeper outperforms other methods at lower levels of $\alpha$.
Figure 7: Performance on VLM classification (left) and captioning tasks (right). Consistent with results in Figures 4 and 6, we see that smaller $\alpha$s lead to improved deferral performance.
@@ 4.2 Decoder-only Setup (Language Models)
We next evaluate Gatekeeper on decoder-only language models. Our setup uses Gemma2B as $\mathcal{M}_{S}$ and Gemma7B as $\mathcal{M}_{L}$ (GemmaTeam et al., 2024), forming a scalable cascade for next-token prediction. As in the classification case, the deferral decision is based solely on the output of $\mathcal{M}_{S}$, using predictive entropy to route uncertain tokens to $\mathcal{M}_{L}$.
Both models are first instruction-tuned on the training split of each dataset to ensure strong base performance and adherence to task format. We then fine-tune $\mathcal{M}_{S}$ with Gatekeeper to reduce its confidence on incorrect token predictions. Evaluation is performed on the validation split using the same metrics as in Section 4.1. We use three representative benchmarks: ARC-e/c (Clark et al., 2018), MMLU (Hendrycks et al., 2020), and GSM8K (Cobbe et al., 2021). These span symbolic reasoning, factual recall, and multi-step arithmetic generation, respectively.
Results. Figure 6 shows that Gatekeeper leads to improved deferral performance and better separation of correct/incorrect predictions, particularly at lower $\alpha$. As with classification, higher $\alpha$ values preserve $\mathcal{M}_{S}$ accuracy but yield weaker separation. In addition to the baseline (Gemma2B with entropy-based deferral), we compare against: (i) the token-level cascading technique from Gupta et al. (2024); and (ii) two uncertainty prompting baselines (see Appendix C.2): Reduce Confidence and Answer “N”. These methods follow Kadavath et al. (2022) and aim to elicit uncertainty via prompting. Overall, we find that Gatekeeper consistently outperforms both prompting and token-level deferral baselines in terms of deferral quality, without modifying the model architecture or inference pipeline.
@@ 4.3 Encoder-Decoder Setup (Vision-Language Models)
We conclude by evaluating Gatekeeper on encoder-decoder vision-language models. Specifically, we use the PaliGemma family (Steiner et al., 2024), which supports tasks like image captioning, VQA, and descriptive classification. $\mathcal{M}_{S}$ is PaliGemma1B, and $\mathcal{M}_{L}$ is PaliGemma7B. Our cascade runs $\mathcal{M}_{S}$ on all inputs and defers to $\mathcal{M}_{L}$ when the predictive entropy of $\mathcal{M}_{S}$ is low. Following Section 4.2, we first fine-tune both models on the task using standard supervised instruction tuning. We then apply Gatekeeper to $\mathcal{M}_{S}$ alone and evaluate the resulting cascade on two classification datasets (VQAv2 (Goyal et al., 2017), AI2D (Hiippala et al., 2021)) and two captioning datasets (Cococap (Lin et al., 2014), Screen2Words (Wang et al., 2021)), covering both closed-form and generative VL tasks.
Factuality Scoring. For classification tasks, we evaluate accuracy and confidence separation as before. For captioning, however, we assess factual alignment between generated and reference captions using the Gemini LLM (GeminiTeam et al., 2023). The model is prompted with: “Are these captions semantically equivalent?” and outputs “Yes” or “No.” We compute the normalized log-likelihood of each response to obtain a factuality score $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$, reflecting the model’s confidence in semantic agreement. Full details are provided in Appendix C.4.
Uncertainty/Factuality Correlation. Since factuality scores are continuous, we adapt our metrics accordingly. Instead of binary accuracy-based separation, we compute the Pearson correlation $\rho\big{(}g_{\text{NENT}}(\mathbf{x}_{i}),s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})\big{)}$ between negative predictive entropy and factuality. We also generalize our deferral performance metric $s_{d}$ to operate on factuality instead of accuracy.
Results. Figure 7 shows our main results. For classification tasks (left), trends match those observed in Sections 4.1 and 4.2: lower $\alpha$ improves deferral but reduces raw accuracy. For captioning tasks (right), Gatekeeper increases the correlation between confidence and factuality, demonstrating effective deferral in generative settings. Prompting baselines from Section 4.2 could not be evaluated: PaliGemma failed to return valid outputs under prompt modifications, likely due to rigid pretraining (Beyer et al., 2024).
@@ 5 Conclusion
Limitations. Despite achieving strong performance across tasks and architectures, several limitations remain: (i) We assume that only $\mathcal{M}_{S}$ is fine-tuned. Although this simplifies deployment and avoids retraining $\mathcal{M}_{L}$, it may overlook gains achievable through joint adaptation. (ii) In language modeling, Gatekeeper may be overly strict: different token sequences can express the same meaning, and penalizing deviations based on exact tokens may suppress valid linguistic variation. Ideally, deferral decisions should reflect semantic correctness rather than surface-level mismatches. (iii) While we evaluate across multiple model families for classification, our experiments in the LLM and VLM settings focus on a single architecture per task, limiting insights into generalization. (iv) Our use of a generative model (Gemini) to score factuality introduces potential noise, as LLMs may produce inconsistent or inaccurate judgments.
Appendix A Broader Impact
This work contributes to the responsible and efficient deployment of machine learning systems by improving the decision-making capabilities of smaller, local models in model cascade architectures. By introducing a loss function that calibrates model confidence with respect to correctness, our approach enhances both the performance and transparency of automated systems that must decide when to act autonomously and when to defer to a more capable model. This design can improve the accessibility and sustainability of machine learning applications by reducing reliance on large, energy-intensive models—particularly important in low-resource environments or edge computing.
At the same time, the ability to fine-tune smaller models to strategically abstain from uncertain predictions raises important considerations for fairness and accountability. In high-stakes applications such as healthcare or finance, improper tuning of the deferral threshold—or uncalibrated confidence estimates—could lead to the systematic denial of service or misallocation of computational resources. Care must be taken to ensure that such systems are thoroughly evaluated not only for average performance but also for differential performance across subgroups. Moreover, the use of large models as fallback decision-makers assumes their correctness, which may not always hold, especially in underrepresented domains. We therefore encourage developers and practitioners to accompany deployments of cascade-based systems with rigorous audits of fairness, reliability, and alignment with human values.
Appendix B Additional Background
B.1 Related Work
B.1.1 LLM Routing
Ding et al. (2024) propose a hybrid LLM inference pipeline that routes each query either to a small on-device model or a larger high-quality model based on the query’s predicted difficulty and a tunable quality threshold. This cost-aware router allows dynamically trading off accuracy for efficiency, enabling up to a 40% reduction in expensive model calls without degrading answer quality. Similarly, Shnitzer et al. (2023) present a method to select the best model from a pool of pre-trained LLMs for each input by learning a “router” on many benchmark tasks. Without requiring labeled examples from the new target task, their approach uses existing datasets to train input-based model selectors, which consistently outperform always using the single best LLM for all queries.
B.1.2 Model Cascade Learning
Nie et al. (2024) introduce an online cascade-learning framework where lightweight models are incrementally trained to imitate a powerful LLM’s decisions on a data stream, deferring to the LLM only when necessary. They cast cascade construction as an imitation-learning problem with theoretical no-regret guarantees, achieving LLM-level accuracy while cutting inference cost by up to 90% and maintaining robustness to distribution shifts over time. Chen et al. (2023) outline strategies for reducing LLM usage cost and present FrugalGPT, a cascade approach that learns to route queries through combinations of smaller or larger LLMs to balance cost and performance. Their experiments show that an adaptive use of multiple models can match the accuracy of the strongest individual LLM (e.g., GPT-4) with up to 98% cost savings. It can also slightly exceed GPT-4’s accuracy at equal cost, highlighting the benefit of cascades that allocate queries to the most appropriate model for each input.
B.1.3 Confidence Calibration in LLMs
Jitkrittum et al. (2023) analyze the classical strategy of confidence-based deferral in model cascades, wherein a model hands off to a stronger model if its confidence is below a threshold, to determine when this simple strategy succeeds or breaks down. They derive the optimal deferral policy in theory and show that naïve confidence thresholds perform well in general but can fail when later models are specialists (only reliable on certain inputs), when there is label noise, or under distribution shift – scenarios where more sophisticated deferral criteria yield better performance. Geng et al. (2023) provide a comprehensive survey of methods for confidence estimation and calibration in LLM outputs. They review recent techniques to quantify uncertainty in large language model predictions, discuss challenges unique to LLMs, and highlight advancements that improve alignment between a model’s reported confidence and its actual accuracy across tasks. Azaria & Mitchell (2023) find evidence that an LLM’s internal activations encode whether or not it is producing a truthful answer, even when the model’s output is incorrect or fabricated. By training a classifier on the model’s hidden state (without fine-tuning the LLM itself), they can often detect when the model is “lying” or unsure, suggesting that large models internally recognize their mistakes or uncertainty despite outwardly confident responses. Similarly, Liu et al. (2024) propose a supervised approach to LLM uncertainty quantification that leverages labeled examples and the model’s hidden representations to predict the correctness of its answers. They show that incorporating features from the model’s internal layers yields significantly improved uncertainty estimates and calibration across diverse tasks, with these gains transferring robustly to new domains. Notably, their method is easy to implement and can be adapted to different levels of model access (black-box vs. white-box), making it widely applicable.
B.1.4 Confidence Verbalization in LLMs
Lin et al. (2022) demonstrate that GPT-3 can be fine-tuned to output a calibrated verbal confidence (e.g., “I’m 90% sure”) along with each answer. This model’s stated confidence levels align well with its true correctness likelihood and remain fairly well-calibrated even under distribution shift, marking the first instance of an LLM explicitly expressing useful uncertainty estimates in natural language. Xiong et al. (2024) thoroughly evaluate black-box methods for eliciting an LLM’s self-reported confidence through prompting and answer sampling. They find that current LLMs tend to verbalize overly high confidence (mirroring human overconfidence), but that carefully designed prompts, consistency checks across multiple sampled answers, and improved aggregation strategies can mitigate this issue. Moreover, larger models generally show better calibration and an improved ability to predict their own failures, though room for further improvement remains in making their expressed uncertainty truly reliable. Mielke et al. (2022) examine whether a conversational agent’s expressed certainty corresponds to its actual knowledge, showing that off-the-shelf dialogue models are poorly “linguistically calibrated.” They demonstrate that a model’s likelihood of giving a correct answer can be estimated via an auxiliary model and used as a control signal to adjust the agent’s responses. The resulting dialogue agent exhibits far less overconfident language when it is likely to be wrong, improving transparency about uncertainty in its answers. Finally, Mahaut et al. (2024) assess the reliability of various methods to estimate an LLM’s factual confidence – the probability that its answer is correct – under both in-domain and paraphrased inputs. Through a rigorous evaluation on QA and fact-checking tasks, they conclude that the most trustworthy confidence scores come from model-introspective approaches (e.g., a trained probe on hidden states), albeit at the cost of requiring full model access and training data. They also highlight that an LLM’s confidence can be unstable under meaning-preserving input variations (paraphrases), underscoring the need for more robust and stable confidence estimation techniques for factual correctness.
B.2 Model Access Levels